OpenStudy (egbeach):

Find the vertices and foci of the hyperbola with equation quantity x minus three squared divided by sixteen minus the quantity of y plus four squared divided by nine = 1

OpenStudy (anonymous):

Is this you problem ? \( \huge \frac{(x -3)^2}{16} - \frac{(y+4)^2}{9} = 1\)

OpenStudy (anonymous):

First find your center. Do you know the center?